OpenStudy (anonymous):

find the sum -7+(-2)

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wat about another one

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

find the sum 4+(-11)

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

one more last one

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

find the difference -8-(-11)

OpenStudy (anonymous):

3

OpenStudy (anonymous):

lets see if i passed

OpenStudy (anonymous):

Okay(:

OpenStudy (anonymous):

yep 100

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

when i need more help ill message u ok

OpenStudy (anonymous):

Anytime

OpenStudy (mathstudent55):

This is how you add two numbers: 1. If both numbers are non-negative, just add them together like you've always been doing. Examples: 2 + 5 = 7; 3 + 0 = 3; 21 + 22 = 43 2. If both numbers are negative, ignore the signs, add the numbers, then the answer is negative. Examples: -4 + (-2) = -6; -32 + (-32) = -64 3. If one number is positive and one number is negative, ignore the negative sign, and subtract the smaller number from the larger number. Then the answer will have the same sign as the larger number. Example 1: -8 + 9 Subtract 9 - 8 = 1 SInce 9 is larger than 8 and 9 is positive, answer is positive, so -8 + 9 =1 Example 2: 18 + (-23) Subtract 23 - 18 = 5 Since 23 is bigger than 18, answer has same sign as -23, so 18 + (-23) = -5